# KCL-PH-TH/2026-03

### v2

@@ Appendix C Parity-odd quadratic-curvature invariants in 4D
where we have assumed a general torsion. Note that $\bar{G}^{-}_{8}$ is the topological Pontryagin term, a total derivative. If we specify our torsion to be the totally antisymmetric tensor $\mathcal{H}_{\mu\nu\lambda}$, which also satisfies the Bianchi identity $\nabla_{[\rho}\mathcal{H}_{\mu\nu\lambda]}=0$, these four parity-odd quadratic invariants are evaluated to be (modulo total derivative terms)
@@
Appendix D Conformal-time analysis of $\lambda$-dependent corrections to the gCS condensate
Going into conformal time, the wave equations for the chiral GW modes (V), (V) become:
@@ Appendix D Conformal-time analysis of $\lambda$-dependent corrections to the gCS condensate
Going into Fourier modes, we get
$\left(1+\lambda\kappa^{2}\left(\frac{b^{\prime}}{a}\right)^{2}\right)h_{L}^{\prime\prime}+\left(2\frac{a^{\prime}}{a}+2\lambda\kappa^{2}\frac{b^{\prime}}{a}\frac{b^{\prime\prime}}{a}\right)h_{L}^{\prime}+k^{2}h_{L}=-\frac{2k\epsilon\kappa^{2}}{a^{2}}\left(b^{\prime\prime}h_{L}^{\prime}+b^{\prime}h_{L}^{\prime\prime}+k^{2}b^{\prime}h_{L}\right)$ (180)
$\left(1+\lambda\kappa^{2}\left(\frac{b^{\prime}}{a}\right)^{2}\right)h_{R}^{\prime\prime}+\left(2\frac{a^{\prime}}{a}+2\lambda\kappa^{2}\frac{b^{\prime}}{a}\frac{b^{\prime\prime}}{a}\right)h_{R}^{\prime}+k^{2}h_{R}=\frac{2k\epsilon\kappa^{2}}{a^{2}}\partial_{z}\left(b^{\prime\prime}h_{R}^{\prime}+b^{\prime}h_{R}^{\prime\prime}+k^{2}b^{\prime}h_{R}\right)$ (181)
or, by bringing all terms to one side, we get:
$\left[1+\lambda\kappa^{2}\left(\frac{b^{\prime}}{a}\right)^{2}+2\epsilon\kappa^{2}\frac{k}{a^{2}}b^{\prime}\right]h_{L}^{\prime\prime}+\left[2\frac{a^{\prime}}{a}+2\lambda\kappa^{2}\frac{b^{\prime}}{a}\frac{b^{\prime\prime}}{a}+2\epsilon\kappa^{2}\frac{k}{a^{2}}b^{\prime\prime}\right]h_{L}^{\prime}+k^{2}\left[1+2\epsilon\kappa^{2}\frac{k}{a^{2}}b^{\prime}\right]h_{L}=0$ (182)
$\left[1+\lambda\kappa^{2}\left(\frac{b^{\prime}}{a}\right)^{2}-2\epsilon\kappa^{2}\frac{k}{a^{2}}b^{\prime}\right]h_{R}^{\prime\prime}+\left[2\frac{a^{\prime}}{a}+2\lambda\kappa^{2}\frac{b^{\prime}}{a}\frac{b^{\prime\prime}}{a}-2\epsilon\kappa^{2}\frac{k}{a^{2}}b^{\prime\prime}\right]h_{R}^{\prime}+k^{2}\left[1-2\epsilon\kappa^{2}\frac{k}{a^{2}}b^{\prime}\right]h_{R}=0$ (183)
@@
$h_{L,R}^{\prime\prime}+\frac{1}{\kappa}P_{L,R}h_{L,R}^{\prime}+k^{2}Q_{L,R}h_{L,R}=0$ (184)
@@
$P_{L,R}$ $=2\frac{\kappa\frac{a^{\prime}}{a}+\lambda\kappa^{3}\frac{b^{\prime}}{a}\frac{b^{\prime\prime}}{a}+l_{L,R}\epsilon\kappa^{3}\frac{k}{a^{2}}b^{\prime\prime}}{1+\lambda\kappa^{2}\left(\frac{b^{\prime}}{a}\right)^{2}+2l_{L,R}\epsilon\kappa^{2}\frac{k}{a^{2}}b^{\prime}}$ (185)
$Q_{L,R}$ $=\frac{1+2l_{L,R}\epsilon\kappa^{2}\frac{k}{a^{2}}b^{\prime}}{1+\lambda\kappa^{2}\left(\frac{b^{\prime}}{a}\right)^{2}+2l_{L,R}\epsilon\kappa^{2}\frac{k}{a^{2}}b^{\prime}}$ (186)
The reader should note that the potentially dangerous modes with momenta i.e. $k$ that lead to the cancellation between
$0\simeq\lambda\kappa^{2}\left(\frac{b^{\prime}}{a}\right)^{2}+2l_{L,R}\epsilon\kappa^{2}\frac{k}{a^{2}}b^{\prime}$
for $l_{L}=-1$ in the denominators of $P_{L},P_{R}$ are superhorizon (frozen) modes ($k<aH$), since they satisfy $k\simeq 0.053\,aH$, hence this condition is never met by the subhorizon modes ($k>aH$) that contribute to the formation of the condensate  [7]. In any case, the difference between $L,R$ modes (birefringence) persists for all subhorizon modes $k$.
Taking into account that $k/a\lesssim\mu=M_{s}$, as appropriate for the UV cutoff of a string-effective theory like the current one (cf. (100)), we can estimate, using (123):
@@
which allows us to approximate $Q$:
@@
where $l_{L}=+1$ and $l_{R}=-1$ and the $\dots$ denote higher-order terms in our small-parameter expansion. Define
$z_{L,R}=e^{\frac{1}{2\kappa}\int^{\eta}P_{L,R}d\tilde{\eta}}$ (190)
such that $\frac{1}{\kappa}P_{L,R}=2\frac{z^{\prime}_{L,R}}{z_{L,R}}$ and the equations become
$h_{L,R}^{\prime\prime}+2\frac{z^{\prime}_{L,R}}{z_{L,R}}h_{L,R}^{\prime}+k^{2}Q_{L,R}h_{L,R}=0$ (191)
@@
$\psi_{L,R}=z_{L,R}h_{L,R},$ (192)
then, $\psi_{L,R}$ satisfies the equation
$\psi_{L,R}^{\prime\prime}+\omega^{2}_{L,R}\psi_{L,R}=0$ (193)
@@
$\omega^{2}_{L,R}=k^{2}Q_{L,R}-\frac{z_{L,R}^{\prime\prime}}{z_{L,R}}$ (194)
@@
$Q_{L,R}=1-\lambda\kappa^{2}\left(\frac{b^{\prime}}{a}\right)^{2}+\mathcal{O}(\lambda\epsilon)$ (195)
@@
$Q_{L,R}=1-\lambda\kappa^{2}\dot{b}^{2}+\mathcal{O}(\lambda\epsilon)$ (196)
@@
$\frac{z_{L,R}^{\prime\prime}}{z_{L,R}}=\frac{P_{L,R}^{2}}{4\kappa^{2}}+\frac{P_{L,R}^{\prime}}{2\kappa}$ (197)
@@
$\frac{z_{L,R}^{\prime\prime}}{z_{L,R}}=\frac{a^{\prime\prime}}{a}+\epsilon\kappa^{2}kl_{L,R}\Delta_{1}+\lambda\kappa^{2}b^{\prime}(\Delta_{1}+\Delta_{2})+\mathcal{O}(\lambda\epsilon)$ (198)
@@
$\frac{z_{L,R}^{\prime\prime}}{z_{L,R}}=a^{2}\Bigg[\left(H^{2}+\frac{\ddot{a}}{a}\right)+\epsilon\kappa^{2}\left(\frac{k}{a}\right)l_{L,R}\left(\dddot{b}+H\ddot{b}-H^{2}\dot{b}-\frac{\ddot{a}}{a}\dot{b}\right)+\lambda\kappa^{2}\dot{b}\left(\dddot{b}+\frac{\ddot{b}^{2}}{\dot{b}}+3H\ddot{b}\right)\Bigg]$ (203)
Therefore, overall, we can evaluate $\omega^{2}_{L,R}=k^{2}Q_{L,R}-\frac{z_{L,R}^{\prime\prime}}{z_{L,R}}$ in cosmic time:
$\begin{split}\omega_{L,R}^{2}=a^{2}\Bigg[\left(\frac{k}{a}\right)^{2}\left(1-\lambda\kappa^{2}\dot{b}^{2}\right)-\left(H^{2}+\frac{\ddot{a}}{a}\right)&-\epsilon\kappa^{2}\left(\frac{k}{a}\right)l_{L,R}\left(\dddot{b}+H\ddot{b}-H^{2}\dot{b}-\frac{\ddot{a}}{a}\dot{b}\right)\\ &-\lambda\kappa^{2}\dot{b}\left(\dddot{b}+\frac{\ddot{b}^{2}}{\dot{b}}+3H\ddot{b}\right)\Bigg]\end{split}$ (204)
This can be rewritten as:
$\omega_{L,R}^{2}=a^{2}\Bigg[\left(\frac{k}{a}\right)^{2}-\left(H^{2}+\frac{\ddot{a}}{a}\right)-\epsilon\kappa^{2}\left(\frac{k}{a}\right)l_{L,R}\left(\dddot{b}+H\ddot{b}-H^{2}\dot{b}-\frac{\ddot{a}}{a}\dot{b}\right)-\lambda\kappa^{2}\dot{b}\left(\dddot{b}+\frac{\ddot{b}^{2}}{\dot{b}}+3H\ddot{b}+\left(\frac{k}{a}\right)^{2}\dot{b}\right)\Bigg]$ (205)
@@
$\psi_{L,R}^{\prime\prime}+\omega^{2}_{L,R}\psi_{L,R}=0\Rightarrow\ddot{\psi}_{L,R}+H\dot{\psi}_{L,R}+\frac{\omega^{2}_{L,R}}{a^{2}}\psi_{L,R}=0$ (206)
@@
$\ddot{\psi}_{L,R}+H\dot{\psi}_{L,R}+\Bigg[\left(\frac{k}{a}\right)^{2}-\left(H^{2}+\frac{\ddot{a}}{a}\right)$ $-\epsilon\kappa^{2}\left(\frac{k}{a}\right)l_{L,R}\left(\dddot{b}+H\ddot{b}-H^{2}\dot{b}-\frac{\ddot{a}}{a}\dot{b}\right)$ (207)
$-\lambda\kappa^{2}\dot{b}\left(\dddot{b}+\frac{\ddot{b}^{2}}{\dot{b}}+3H\ddot{b}+\left(\frac{k}{a}\right)^{2}\dot{b}\right)\Bigg]\,\psi_{L,R}=0\,.$ (208)
For $\dot{b}\simeq{\rm const}$ and $\frac{\ddot{a}}{a}\simeq H^{2}$ we have that
$\ddot{\psi}_{L,R}+H\dot{\psi}_{L,R}+\Bigg[\left(\frac{k}{a}\right)^{2}-2H^{2}+2\epsilon\kappa^{2}\left(\frac{k}{a}\right)l_{L,R}H^{2}\dot{b}-\lambda\kappa^{2}\left(\frac{k}{a}\right)^{2}\dot{b}^{2}\Bigg]\,\psi_{L,R}=0\,.$ (209)
$\ddot{\psi}_{L,R}+H\dot{\psi}_{L,R}+\Bigg[\left(\frac{k}{a}\right)^{2}\left(1-\lambda\kappa^{2}\dot{b}^{2}\right)-2H^{2}+2\epsilon\kappa^{2}\left(\frac{k}{a}\right)l_{L,R}H^{2}\dot{b}\Bigg]\,\psi_{L,R}=0\,.$ (210)
We rescale $\frac{1}{\tilde{a}}=\frac{\sqrt{1-\lambda\kappa^{2}\dot{b}^{2}}}{a}$ (cf. (131)) and get:
$\ddot{\psi}_{L,R}+H\dot{\psi}_{L,R}+\Bigg[\left(\frac{k}{\tilde{a}}\right)^{2}-2H^{2}+2\epsilon\kappa^{2}\frac{1}{\sqrt{1-\lambda\kappa^{2}\dot{b}^{2}}}\left(\frac{k}{\tilde{a}}\right)l_{L,R}H^{2}\dot{b}\Bigg]\,\psi_{L,R}=0\,.$ (211)
$\ddot{\psi}_{L,R}+H\dot{\psi}_{L,R}+\Bigg[\left(\frac{k}{\tilde{a}}\right)^{2}-2H^{2}+2\epsilon\kappa^{2}\left(\frac{k}{\tilde{a}}\right)l_{L,R}H^{2}\dot{b}+\mathcal{O}(\epsilon\lambda)\Bigg]\,\psi_{L,R}=0\,.$ (212)
This is nothing other than the corresponding wave equations of chiral GW linear polarizations of [7], expressed in terms of the rescaled scale factor $a^{\lambda}$ (131). Following the canonical quantization approach of that work, then, in the evaluation of the gravitational CS condensate for weak $\dot{b}$, leads to (135) in the text.

### v3

@@ Appendix C Parity-Odd Quadratic Curvature Invariants in 4D
where we assume a general torsion. Note that $\bar{G}^{-}_{8}$ is the topological Pontryagin term, a total derivative. If we specify our torsion to be the totally antisymmetric tensor $\mathcal{H}_{\mu\nu\lambda}$, which also satisfies the Bianchi identity $\nabla_{[\rho}\mathcal{H}_{\mu\nu\lambda]}=0$, these four parity-odd quadratic invariants are evaluated to be modulo total derivative terms as follows:
@@
Appendix D Conformal Time Analysis of $\lambda$-Dependent Corrections to the gCS Condensate
Going into conformal time, the wave equations for the chiral GW modes (V) and (V) become
@@ Appendix D Conformal Time Analysis of $\lambda$-Dependent Corrections to the gCS Condensate
Going into Fourier modes:
$h_{L,R}(\eta,\vec{x})=\int\frac{d^{3}k}{(2\pi)^{\frac{3}{2}}}e^{i\vec{k}\cdot\vec{x}}h_{L,R,\vec{k}}(\eta),$ (183)
and substituting, we get
$\left(1+\lambda\kappa^{2}\left(\frac{b^{\prime}}{a}\right)^{2}\right)h_{L,\vec{k}}^{\prime\prime}+\left(2\frac{a^{\prime}}{a}+2\lambda\kappa^{2}\frac{b^{\prime}}{a}\frac{b^{\prime\prime}}{a}\right)h_{L,\vec{k}}^{\prime}+k^{2}h_{L,\vec{k}}$
$=-\frac{2k\epsilon\kappa^{2}}{a^{2}}l_{\vec{k}}\left(b^{\prime\prime}h_{L,\vec{k}}^{\prime}+b^{\prime}h_{L,\vec{k}}^{\prime\prime}+k^{2}b^{\prime}h_{L,\vec{k}}\right)$ (184)
$\left(1+\lambda\kappa^{2}\left(\frac{b^{\prime}}{a}\right)^{2}\right)h_{R,\vec{k}}^{\prime\prime}+\left(2\frac{a^{\prime}}{a}+2\lambda\kappa^{2}\frac{b^{\prime}}{a}\frac{b^{\prime\prime}}{a}\right)h_{R,\vec{k}}^{\prime}+k^{2}h_{R,\vec{k}}$
$=\frac{2k\epsilon\kappa^{2}}{a^{2}}l_{\vec{k}}\left(b^{\prime\prime}h_{R,\vec{k}}^{\prime}+b^{\prime}h_{R,\vec{k}}^{\prime\prime}+k^{2}b^{\prime}h_{R,\vec{k}}\right)$ (185)
where $l_{\vec{k}}=+1$ and $l_{-\vec{k}}=-1$. By bringing all terms to one side, we get
$\begin{split}\left[1+\lambda\kappa^{2}\left(\frac{b^{\prime}}{a}\right)^{2}+2\epsilon\kappa^{2}l_{\vec{k}}\frac{k}{a^{2}}b^{\prime}\right]h_{L,\vec{k}}^{\prime\prime}&+\left[2\frac{a^{\prime}}{a}+2\lambda\kappa^{2}\frac{b^{\prime}}{a}\frac{b^{\prime\prime}}{a}+2\epsilon\kappa^{2}l_{\vec{k}}\frac{k}{a^{2}}b^{\prime\prime}\right]h_{L,\vec{k}}^{\prime}\\ &+k^{2}\left[1+2\epsilon\kappa^{2}l_{\vec{k}}\frac{k}{a^{2}}b^{\prime}\right]h_{L,\vec{k}}=0\end{split}$ (186)
$\begin{split}\left[1+\lambda\kappa^{2}\left(\frac{b^{\prime}}{a}\right)^{2}-2\epsilon\kappa^{2}l_{\vec{k}}\frac{k}{a^{2}}b^{\prime}\right]h_{R,\vec{k}}^{\prime\prime}&+\left[2\frac{a^{\prime}}{a}+2\lambda\kappa^{2}\frac{b^{\prime}}{a}\frac{b^{\prime\prime}}{a}-2\epsilon\kappa^{2}l_{\vec{k}}\frac{k}{a^{2}}b^{\prime\prime}\right]h_{R,\vec{k}}^{\prime}\\ &+k^{2}\left[1-2\epsilon\kappa^{2}l_{\vec{k}}\frac{k}{a^{2}}b^{\prime}\right]h_{R,\vec{k}}=0\end{split}$ (187)
@@
$h_{L,R,\vec{k}}^{\prime\prime}+\frac{1}{\kappa}P_{L,R,\vec{k}}h_{L,R,\vec{k}}^{\prime}+k^{2}Q_{L,R,\vec{k}}h_{L,R,\vec{k}}=0$ (188)
@@
$P_{L,R,\vec{k}}$ $=2\frac{\kappa\frac{a^{\prime}}{a}+\lambda\kappa^{3}\frac{b^{\prime}}{a}\frac{b^{\prime\prime}}{a}+l_{L,R}\epsilon\kappa^{3}l_{\vec{k}}\frac{k}{a^{2}}b^{\prime\prime}}{1+\lambda\kappa^{2}\left(\frac{b^{\prime}}{a}\right)^{2}+2l_{\vec{k}}\,l_{L,R}\epsilon\kappa^{2}\frac{k}{a^{2}}b^{\prime}}$ (189)
$Q_{L,R,\vec{k}}$ $=\frac{1+2l_{\vec{k}}\,l_{L,R}\epsilon\kappa^{2}\frac{k}{a^{2}}b^{\prime}}{1+\lambda\kappa^{2}\left(\frac{b^{\prime}}{a}\right)^{2}+2l_{\vec{k}}\,l_{L,R}\epsilon\kappa^{2}\frac{k}{a^{2}}b^{\prime}}$ (190)
The reader should note that the potentially dangerous modes with momenta, i.e., $k$, that lead to the cancellation of
$0\simeq\lambda\kappa^{2}\left(\frac{b^{\prime}}{a}\right)^{2}+2l_{\vec{k}}\,l_{L,R}\epsilon\kappa^{2}\frac{k}{a^{2}}b^{\prime}$
for $l_{\vec{k}}l_{L,R}=-1$ in the denominators of $P_{L,\vec{k}},P_{R,\vec{k}}$ are superhorizon (frozen) modes ($k<aH$), since they satisfy $k\simeq 0.053\,a~H$; hence, this condition is never met by the subhorizon modes ($k>aH$) that contribute to the formation of the condensate [7]. In any case, the difference between $L,R$ modes (birefringence) persists for all subhorizon modes $k$.
Taking into account that $k/a\lesssim\mu=M_{s}$, as is appropriate for the UV cutoff of a string-effective theory like the current one (cf. (100)), we can estimate, using (124), the following:
@@
which allows us to approximate $Q$ as follows:
@@
where $l_{L}=+1$ and $l_{R}=-1$ and the $\dots$ denotes higher-order terms in our small-parameter expansion. Define
$z_{L,R}=e^{\frac{1}{2\kappa}\int^{\eta}P_{L,R,\vec{k}}d\tilde{\eta}}$ (194)
such that $\frac{1}{\kappa}P_{L,R,\vec{k}}=2\frac{z^{\prime}_{L,R,\vec{k}}}{z_{L,R,\vec{k}}}$ and the equations become
$h_{L,R,\vec{k}}^{\prime\prime}+2\frac{z^{\prime}_{L,R,\vec{k}}}{z_{L,R,\vec{k}}}h_{L,R,\vec{k}}^{\prime}+k^{2}Q_{L,R,\vec{k}}h_{L,R,\vec{k}}=0$ (195)
@@
$\psi_{L,R,\vec{k}}=z_{L,R,\vec{k}}h_{L,R,\vec{k}},$ (196)
then, $\psi_{L,R,\vec{k}}$ satisfies the equation
$\psi_{L,R,\vec{k}}^{\prime\prime}+\omega^{2}_{L,R,\vec{k}}\psi_{L,R,\vec{k}}=0$ (197)
@@
$\omega^{2}_{L,R,\vec{k}}=k^{2}Q_{L,R,\vec{k}}-\frac{z_{L,R,\vec{k}}^{\prime\prime}}{z_{L,R,\vec{k}}}$ (198)
@@
$Q_{L,R,\vec{k}}=1-\lambda\kappa^{2}\left(\frac{b^{\prime}}{a}\right)^{2}+\mathcal{O}(\lambda\epsilon)$ (199)
@@
$Q_{L,R,\vec{k}}=1-\lambda\kappa^{2}\dot{b}^{2}+\mathcal{O}(\lambda\epsilon)$ (200)
@@
$\frac{z_{L,R,\vec{k}}^{\prime\prime}}{z_{L,R,\vec{k}}}=\frac{P_{L,R,\vec{k}}^{2}}{4\kappa^{2}}+\frac{P_{L,R,\vec{k}}^{\prime}}{2\kappa}$ (201)
@@
$\frac{z_{L,R,\vec{k}}^{\prime\prime}}{z_{L,R,\vec{k}}}=\frac{a^{\prime\prime}}{a}+\epsilon\kappa^{2}kl_{\vec{k}}\,l_{L,R}\Delta_{1}+\lambda\kappa^{2}b^{\prime}(\Delta_{1}+\Delta_{2})+\mathcal{O}(\lambda\epsilon)$ (202)
@@
$\frac{z_{L,R,\vec{k}}^{\prime\prime}}{z_{L,R,\vec{k}}}=a^{2}\Bigg[\left(H^{2}+\frac{\ddot{a}}{a}\right)+\epsilon\kappa^{2}\left(\frac{k}{a}\right)l_{\vec{k}}\,l_{L,R}\left(\dddot{b}+H\ddot{b}-H^{2}\dot{b}-\frac{\ddot{a}}{a}\dot{b}\right)+\lambda\kappa^{2}\dot{b}\left(\dddot{b}+\frac{\ddot{b}^{2}}{\dot{b}}+3H\ddot{b}\right)\Bigg]$ (207)
Therefore, overall, we can evaluate $\omega^{2}_{L,R,\vec{k}}=k^{2}Q_{L,R,\vec{k}}-\frac{z_{L,R,\vec{k}}^{\prime\prime}}{z_{L,R,\vec{k}}}$ in cosmic time:
$\begin{split}\omega_{L,R,\vec{k}}^{2}=a^{2}\Bigg[\left(\frac{k}{a}\right)^{2}\left(1-\lambda\kappa^{2}\dot{b}^{2}\right)-\left(H^{2}+\frac{\ddot{a}}{a}\right)&-\epsilon\kappa^{2}\left(\frac{k}{a}\right)l_{\vec{k}}\,l_{L,R}\left(\dddot{b}+H\ddot{b}-H^{2}\dot{b}-\frac{\ddot{a}}{a}\dot{b}\right)\\ &-\lambda\kappa^{2}\dot{b}\left(\dddot{b}+\frac{\ddot{b}^{2}}{\dot{b}}+3H\ddot{b}\right)\Bigg]\end{split}$ (208)
This can be rewritten as
$\begin{split}\omega_{L,R,\vec{k}}^{2}=a^{2}\Bigg[\left(\frac{k}{a}\right)^{2}-\left(H^{2}+\frac{\ddot{a}}{a}\right)&-\epsilon\kappa^{2}\left(\frac{k}{a}\right)l_{\vec{k}}l_{L,R}\left(\dddot{b}+H\ddot{b}-H^{2}\dot{b}-\frac{\ddot{a}}{a}\dot{b}\right)\\ &-\lambda\kappa^{2}\dot{b}\left(\dddot{b}+\frac{\ddot{b}^{2}}{\dot{b}}+3H\ddot{b}+\left(\frac{k}{a}\right)^{2}\dot{b}\right)\Bigg]\end{split}$ (209)
@@
$\psi_{L,R,\vec{k}}^{\prime\prime}+\omega^{2}_{L,R,\vec{k}}\psi_{L,R,\vec{k}}=0\Rightarrow\ddot{\psi}_{L,R,\vec{k}}+H\dot{\psi}_{L,R,\vec{k}}+\frac{\omega^{2}_{L,R,\vec{k}}}{a^{2}}\psi_{L,R,\vec{k}}=0$ (210)
@@
$\ddot{\psi}_{L,R,\vec{k}}+H\dot{\psi}_{L,R,\vec{k}}+\Bigg[\left(\frac{k}{a}\right)^{2}-\left(H^{2}+\frac{\ddot{a}}{a}\right)$ $-\epsilon\kappa^{2}\left(\frac{k}{a}\right)l_{\vec{k}}\,l_{L,R}\left(\dddot{b}+H\ddot{b}-H^{2}\dot{b}-\frac{\ddot{a}}{a}\dot{b}\right)$
$-\lambda\kappa^{2}\dot{b}\left(\dddot{b}+\frac{\ddot{b}^{2}}{\dot{b}}+3H\ddot{b}+\left(\frac{k}{a}\right)^{2}\dot{b}\right)\Bigg]\,\psi_{L,R,\vec{k}}=0\,.$ (211)
For $\dot{b}\simeq{\rm const}$ and $\frac{\ddot{a}}{a}\simeq H^{2}$, we have that
$\ddot{\psi}_{L,R,\vec{k}}+H\dot{\psi}_{L,R,\vec{k}}+\Bigg[\left(\frac{k}{a}\right)^{2}-2H^{2}+2\epsilon\kappa^{2}\left(\frac{k}{a}\right)l_{\vec{k}}\,l_{L,R}H^{2}\dot{b}-\lambda\kappa^{2}\left(\frac{k}{a}\right)^{2}\dot{b}^{2}\Bigg]\,\psi_{L,R,\vec{k}}=0\,.$ (212)
or, equivalently
$\ddot{\psi}_{L,R,\vec{k}}+H\dot{\psi}_{L,R,\vec{k}}+\Bigg[\left(\frac{k}{a}\right)^{2}\left(1-\lambda\kappa^{2}\dot{b}^{2}\right)-2H^{2}+2\epsilon\kappa^{2}\left(\frac{k}{a}\right)l_{\vec{k}}\,l_{L,R}H^{2}\dot{b}\Bigg]\,\psi_{L,R,\vec{k}}=0\,.$ (213)
We rescale $\frac{1}{\tilde{a}}=\frac{\sqrt{1-\lambda\kappa^{2}\dot{b}^{2}}}{a}$ (cf. (132)) and get
$\ddot{\psi}_{L,R,\vec{k}}+H\dot{\psi}_{L,R,\vec{k}}+\Bigg[\left(\frac{k}{\tilde{a}}\right)^{2}-2H^{2}+2\epsilon\kappa^{2}\frac{1}{\sqrt{1-\lambda\kappa^{2}\dot{b}^{2}}}\left(\frac{k}{\tilde{a}}\right)l_{\vec{k}}\,l_{L,R}H^{2}\dot{b}\Bigg]\,\psi_{L,R,\vec{k}}=0\,$, (214)
which, upon expanding, becomes
$\ddot{\psi}_{L,R,\vec{k}}+H\dot{\psi}_{L,R,\vec{k}}+\Bigg[\left(\frac{k}{\tilde{a}}\right)^{2}-2H^{2}+2\epsilon\kappa^{2}\left(\frac{k}{\tilde{a}}\right)l_{\vec{k}}\,l_{L,R}H^{2}\dot{b}+\mathcal{O}(\epsilon\lambda)\Bigg]\,\psi_{L,R,\vec{k}}=0\,.$ (215)
This is nothing other than the corresponding wave equations of chiral GW linear polarizations of [7], expressed in terms of the rescaled scale factor $a^{\lambda}$ (132). Following the canonical quantization approach of that work, then, in the evaluation of the gravitational CS condensate for a weak $\dot{b}$, leads to (136) in the text.